OpenStudy (anonymous):

If f(x) = |x – 5| + 2, find f(3).

OpenStudy (zzr0ck3r):

what is 3-5?

OpenStudy (anonymous):

2

OpenStudy (zzr0ck3r):

you sure?

OpenStudy (zzr0ck3r):

three minus five

OpenStudy (anonymous):

5-3 is 2. you cant do 3-5

OpenStudy (zzr0ck3r):

sure you can, \(|3-5|+2 = |-2|+2=2+2=4\)

OpenStudy (zzr0ck3r):

it just so happens that \(|3-5|=|5-3| \text{ but } 3-5 \ne 5-3\)

OpenStudy (zzr0ck3r):

and this is the power of the absolute value:)